please answer noth questions and circle correct answer
In September, Larson Inc. sold 40,000 units of its only product for $240,000 and incurred a total cost of $225,000, of which $25,000 is fixed costs. The flexible budget for September showed total sales of $300,000. Among variances of the period were: total variable cost flexible-budget variance, $8,000U; total flexible-budget variance, $63,000U and, sales volume variance, in terms of contribution margin, $27,00oU. The total number of budgeted units reflected in the master budget for September was O 36,000 units. O 40,000 units O45,000 units O48,000 units. O50,000 units
Barstow Manufacturing Company has two service departments-product design and engineering support, and two The distribution ion of each service department's efforts to the other departments is shown below FROM то Design Support Assembly Finishing Design 0% 10% Support 20% 30% 45% 60% 35% 0% The direct operating costs of the departments (including both variable and fixed costs) were as foll Design Engineering Support Assembly Finishing $140,000 $160,000 $550,000 $B40,000 The total cost accumulated in the assembly department using the step method is (calculate all ratios and percentages to 4 decimal places, for esamoie a s and rouod a dollar amounts to the nearest whole dollar; assume that the design department goes first $1,000,125 $1,890,000 o $689,875. $642,000. $923,000
Answer #1

Ans-1-

Actual    Flexible budget       Master
Units 40000 40000 50000 Units
Sales 240000 60000 U 300000 75000 U 375000
Variable cost 200000 8000 U 192000 48000 F 240000
Contribution margin 40000 68000 U 108000 27000 U 135000
Fixed cost 25000 5000 F 30000 0 30000
Net income 15000 63000 U 78000 27000 U 105000

So total sales revenue = 135000*300000/108000 = 375000

Sales price per unit = 300000/40000= 7.5

As per master budget sales units = 375000/7.5 = 50000 units

Hence, The correct option is E.-50,000 units

2-

Design Support Assembly Finishing
Design 0% 10% 30% 60%
Support 20% 0% 45% 35%
Direct Cost Design Support Assembly Finishing
Base Cost 140,000 160,000 550,000 840,000
Design 140,000

14,000

(140,000*10%)

42,000

(140,000*30%)

84,000

(140,000*60%)

Engineering support

174,000

(160,000+14,000)

97,875

(174,000*45%/ (45%+35%)

76,125

(174,000*35%/ (45%+35%)

689,875

Total cost accumulated to assembly department is $689,875

Hence, the correct option is C-689,875